OpenStudy (anonymous):

medal! please help Find the missing term of the geometric sequence. 45, ____, 1620. A) 51 B) 6 C) 270 D) 9720

OpenStudy (zzr0ck3r):

\(\frac{x}{45}=\frac{1620}{x}\)

OpenStudy (zzr0ck3r):

Solve for x

OpenStudy (solomonzelman):

45x²=1620

OpenStudy (zzr0ck3r):

x^2=1620*45 @SolomonZelman

OpenStudy (solomonzelman):

You will have 2 means.

OpenStudy (solomonzelman):

You options are INCOMPLETE.

OpenStudy (zzr0ck3r):

@SolomonZelman x^2=1620*45 not 45x^2=1620

OpenStudy (zzr0ck3r):

the option is there...

OpenStudy (solomonzelman):

Nope

OpenStudy (solomonzelman):

Look closer, you go ` r² ` from 45 to get 1620

OpenStudy (solomonzelman):

So the common ratio can be -6 or 6.

OpenStudy (zzr0ck3r):

if a,b,c is geometric b/a = c/b so x/45=1620/x ie x^2=1620*45

OpenStudy (solomonzelman):

you don't multiply them, it makes no sense 45 × r = missing mean 45 × r × r = 1620 45 r² = 1620 And then divide both sides by 45.

OpenStudy (zzr0ck3r):

do you agree that if we have geometric series a,b,c,d... then b/a=c/d=d/c?

OpenStudy (zzr0ck3r):

if you agree with that then you must agree with x/45 = 1620/x then you also must agree that x^2=45*1620 then you must agree that x = 270

OpenStudy (solomonzelman):

b/a = d/c (not c/d)

OpenStudy (zzr0ck3r):

im doing b/a = c/b here b = x a = 45 , c = 1620

OpenStudy (zzr0ck3r):

mine should say b/a=c/b=d/c?

OpenStudy (solomonzelman):

x × 45 = 1620 ÷ x Like this

OpenStudy (zzr0ck3r):

ok man...

OpenStudy (solomonzelman):

where each side of the equation is equivalent to the missing mean

OpenStudy (solomonzelman):

Or am I wrong ?

OpenStudy (zzr0ck3r):

again for geometric sequence a,b,c we MUST have b/a=c/b here we have 45,x,1620 so x/45=1620/x

OpenStudy (zzr0ck3r):

I think I have made it clear that I think your wrong:P

OpenStudy (zzr0ck3r):

or give me your answer and ill show its wrong:P

OpenStudy (zzr0ck3r):

the one that is not an option that is.

OpenStudy (solomonzelman):

My solution comes up with option C

OpenStudy (zzr0ck3r):

45x^2=1620 x^2=1620/45 gives x = 6 is what you said before

OpenStudy (solomonzelman):

I might be not good at math, but this I know for sure, \(\huge\color{blue}{ t_{n}\times r=\frac{t_{n+2}}{r} }\)

OpenStudy (zzr0ck3r):

if you are getting x = 720 then you changed something

OpenStudy (solomonzelman):

No, 36 is r²

OpenStudy (solomonzelman):

and r=6

OpenStudy (solomonzelman):

or -6

OpenStudy (zzr0ck3r):

lol I solved for x, the missing thing. you solved for r

OpenStudy (zzr0ck3r):

then you will find r to get the missing thing

OpenStudy (zzr0ck3r):

:)

OpenStudy (solomonzelman):

well, we are both solving for the common ratio.

OpenStudy (zzr0ck3r):

you get 6 from x^2=1620/45 I get 270 from x^2=1620*45

OpenStudy (zzr0ck3r):

we both need to define our variables:)

OpenStudy (solomonzelman):

I get x=6 from x²=1620 / 45

OpenStudy (solomonzelman):

and what makes no sense about what get ?

OpenStudy (solomonzelman):

All I am saying is that the common ratio is 6 or -6, and there are 2 geometric means -270 or 270. And the best choice is C, even though it is not complete, since it is missing another mean of -270.

OpenStudy (zzr0ck3r):

yes, again you solved for the ratio, I solved for the missing term. They should have put a 4th term on there to show that it would either be positive or negative.